OpenStudy (anonymous):

Order the group of quadratic functions from widest to narrowest. y=3x^2, y=5x^2, y=7x^2

OpenStudy (anonymous):

i think its just backwards, i think it will be y=7x^2, y=5x^2, y=3x^2

OpenStudy (anonymous):

You think im correct? This is a test to i gotta be sure its right. You kno:0

OpenStudy (anonymous):

@HMT @rose21 what u guys think?

OpenStudy (anonymous):

Girl, math is not my best subject. I suggest using www.mathway.com. That's what I use when I can't figure out a problem!

OpenStudy (anonymous):

i use it too, but see i needa pput these from widest to narrowest lol so i cant do it on there u kno

OpenStudy (anonymous):

But thank u:)

OpenStudy (anonymous):

Oh, lol. Yea then I can't do much help for you, sorry. No problem :)

OpenStudy (anonymous):

@nirmalnema can u help

OpenStudy (nirmalnema):

y=7x^2>y=5x^2> y=3x^2

OpenStudy (anonymous):

So i was right?:)

OpenStudy (nirmalnema):

yup

OpenStudy (anonymous):

Yupp , you were correct . You so smart lol

OpenStudy (anonymous):

lol No No:) i hate math

OpenStudy (anonymous):

I hate it too , but you have to admir you are smart .

OpenStudy (anonymous):

lol:) why thanks, made my day

OpenStudy (anonymous):

your welcome ( :